OpenStudy (anonymous):

geometry: The perimeter of a driveway is 168 ft. The length is 12 ft. longer than three times the width. What are the dimensions of the driveway?

OpenStudy (anonymous):

L=3W+12 2L+2W=168 These are statements made to you. Use the information from the first statement to solve for the other variable in the second statement. Doing this you get: 2W+2(3W+12)=168 2W+6W+24=168 8W=144 W=18 The width is 18. Good. What is the length? Use one of your true statements to solve for the other variable. L=3(18)+12 L=54+12 L=66 To conclude, The Width is 18 and the Length is 66.

OpenStudy (anonymous):

ty so much

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yes